# TACO: Enhancing Multimodal In-context Learning via Task Mapping-Guided Sequence Configuration

### v1

@@ E.2 Training Data Construction Details
To determine the optimal $n$-th demonstration $x_{n}$ for a sequence $S_{i}^{n-1}$ with $n-1$ ICDs, we select the candidate that maximizes the incremental gain in $\mathcal{C}_{\mathcal{M}}$:
$x_{n}=\underset{x\in D_{i}}{argmax}[\mathcal{S}_{\mathcal{C}_{\mathcal{M}}}(S^{n-1}_{i}+x)-\mathcal{S}_{\mathcal{C}_{\mathcal{M}}}\mathcal{M}(S^{n-1}_{i})].$ (31)
(3). We employ beam search with a beam size of $2N$, ensuring that for each $\hat{x}$, the top $2N$ optimal sequences are included in $D_{S}$. As a result, the final sequence set $D_{S}$ consists of $2N\times k$ $N$-shot sequences, providing refined training data for the model.
@@ E.3 Baselines
Various baseline methods are used to evaluate the model’s performance, ranging from random sample to different SOTA retrieval strategies. The following is a description of the baselines used in our experiments.
@@
3. ImageQuery2ImageQuery (IQ2IQ): During the retrieval process, both the image embeddings $I_{i}$ and the query embeddings $Q_{i}$ of each demonstration $(I_{i},Q_{i},R_{i}$ are used. These embeddings are compared to the embedding of the concatenated query sample $(\hat{I},\hat{Q})$ and the retrieval is based on the joint similarity between the images and the queries.
@@
5. DEmO: DEmO is a two-stage, data-free framework for configuring an optimal in-context sequence using influence, a concept that has become increasingly popular in ICD selection. In the first stage, it draws $N$ random permutations $\{\pi_{i}\}_{i=1}^{N}$ of the candidate support set and measures each permutation’s label-fairness by computing its content-free entropy:
@@
where $C_{\pi}$ is the prompt constructed by $\pi$ with a content-free token in place of the query. The top-$K$ permutations with the highest $E(\pi)$ are retained as the candidate set $\Pi$.
@@
Finally, permutation $\pi^{*}=\arg\max_{\pi\in\Pi}I(x_{t};\pi)$ is chosen for the actual prediction. This targeted re-ranking ensures that each test sample uses the example order most “influential” to its correct classification, without relying on any additional labeled data.
@@ E.4 Results and Analysis
VQA Captioning Classification Hybrid Fast CLEVR
@@
We can go deep into the per-model results in Tabel 8. The findings are as follows: (1) TACO exhibits the best performance in all but three tasks across nine datasets and five LVLMs, demonstrating its great efficiency and generalization. Upon examining the outputs, we observe that GPT-4V tends to deviate from the ICD format and produce redundant information more easily than open-source LVLMs, aligning with (Wu et al., 2023a). This results in the quality improvement of the ICL sequence not always translating into stable ICL performance gains for GPT-4V, which may explain why TACO did not achieve the best performance in two of its tasks. (2) For tasks like VizWiz and Hybrid, TACO consistently improves the quality of sequence generation in all LVLMs compared to similarity-based models, demonstrating the importance of increasing task semantics for complex task mappings. We find that the performance gains from TACO are not directly related to the model’s intrinsic ability on these tasks. Unlike simpler tasks like classification, for tasks with complex mappings, task semantics still has a significant impact, even when LVLMs exhibit strong few-shot learning abilities. This shows that models with strong ICL capabilities on certain tasks retain, and even strengthen, their ability to leverage task semantics, underscoring the value of improving ICL sequence quality.
@@
Efficiency analyses. ICL is widely adopted for its efficiency; therefore, efficiency was a primary focus in the design of TACO. TACO exhibits high efficiency during both training and inference. Firstly, TACO is a lightweight language model composed solely of a fusion module and four transformer decoder blocks. With only 140M parameters, its training cost is extremely low compared to LLMs. Moreover, owing to its specialized training objectives, TACO prioritizes data quality over sheer quantity, enabling effective training with a relatively small amount of high-quality data. Simultaneously, our approach to constructing training data is highly efficient. By leveraging LVLM for self-assessment, we significantly reduce the overhead associated with incorporating external metrics. For instance, when using CIDEr to construct training data for the image captioning task, the costs are nearly 9 times higher than those of our current method. To further validate TACO’s training efficiency, we compare its training time with that of 4-layer Lever-LM on the same training sets.
@@ Appendix F Additional Ablation Study
F.1 Training Data
In constructing our training sequences, we use Oracle that approximates an LVLM’s optimal multimodal ICL. Training TACO on sequences generated by Oracle enables it to learn how to configure ICL inputs with correct task mappings, especially at the global level across the demonstration set. To test the robustness of this learning process, we replace Oracle with similarity-based methods and examine how this change affects TACO’s accuracy in inferring task mappings.
Method VQAv2 VizWiz OK-VQA Flickr30K MSCOCO Hatefulmemes Hybrid
TACO(RS) 62.38 47.69 54.47 98.31 115.83 76.49 38.62
TACO(I2I) 61.95 47.28 53.86 98.74 117.25 77.46 36.90
TACO(IQ2IQ) 64.37 50.18 59.23 99.17 118.68 78.93 41.05
TACO(Oracle) 66.75 52.07 61.54 99.62 119.47 80.59 45.22
As TACO is training-based, one of the most important aspects is high-quality data. Table 13 demonstrates that using Oracle as the method to construct the training data is optimal. Since our approach leverages TACO to capture how LVLMs understand and utilize task mapping, the training data that best reflects the internal mechanisms of the LVLM is most effective. Moreover, it can be observed that TACO, when combined with RS, I2I, and IQ2IQ, brings significant performance improvements over using RS, I2I, and IQ2IQ in isolation. This indicates that our method can mitigate the inherent limitations of retrieval strategies through training, thereby further enhancing the practical value of TACO.
F.2 Input Embeddings
Ta VQAv2 MSCOCO Hatefulmemes Hybrid Fast CLEVR
@@ F.2 Input Embeddings
where $f_{I},f_{Q}$ and $f_{R}$ denote the weights computed using a softmax function applied the linear transformations, ensuring their sum equals 1. Additionally, we apply regularization to the weights: $f_{I}^{2}+f_{Q}^{2}+f_{R}^{2}\leq\theta$ to prevent excessive reliance on specific components.
@@ F.3 Instruction
Table 15: Results of TACO with diverse instruction types. The highest scores are highlighted in bold. $Inst^{\prime}\to Inst$ means using $Inst$ during the initialization of $TG$.
$Inst^{\prime}$ $Inst$ VQAv2 VizWiz OKVQA Hybrid
@@
Table 16: Results of TACO under various $Inst^{\prime}$-$Inst$ combinations. $Inst^{\prime}$ represents the style used for initializing $TG$, while $Inst$ refers to the style actually incorporated into the prompt.
F.4 Generalization Test
@@
For NLP evaluation, we utilize the Rule Learning part of the latest benchmark, ICLEval. ICLEval is designed to assess the ICL abilities of LLMs, focusing on two main sub-abilities: exact copying and rule learning. The Rule Learning part evaluates how well LLMs can derive and apply rules from examples in the context. This includes tasks such as format learning, where models must replicate and adapt formats from given examples, and order and statistics-based rule learning, where the model must discern and implement patterns such as item sequencing or handling duplications. These tasks challenge LLMs to go beyond language fluency, testing their ability to generalize from context in diverse scenarios. Examples of $(Q,R)$ pairs can be found in Table 19. For all tasks, we use exact match scores to evaluate the predictions with the labels.
@@
F.5 Revisiting Task Mapping Framework
@@
Besides evaluating performance, we also employ in-context lens to examine the average evolution of the internal outputs in LVLM by these methods under HM settings. Figure 10 illustrates the evolution of the LVLM’s internal reasoning as it learns ICL sequences generated by different methods. The results show that sequences produced by TACO enable the model to most effectively infer task mappings and generate accurate responses. These findings further validate TACO’s effectiveness and confirm the robustness of our task mapping framework.

### v2

@@ E.2 Training Data Construction Details
To determine the optimal $n$-th demonstration $x_{n}$ for a sequence $S_{i}^{n-1}$ with n1n-1 ICDs, we select the candidate that maximizes the incremental gain in $\mathcal{C}_{\mathcal{M}}$:
$x_{n}=\underset{x\in D_{i}}{argmax}[{\mathcal{C}_{\mathcal{M}}}(S^{n-1}_{i}+x)-{\mathcal{C}_{\mathcal{M}}}(S^{n-1}_{i})].$ (29)
(3). We employ beam search with a beam size of $2N$, ensuring that for each $\hat{x}$, the top $2N$ optimal sequences are included in $D_{S}$. As a result, the final sequence set $D_{S}$ consists of $2N\times k$ NN-shot sequences, providing refined training data for the model.
@@ E.3 Baselines
Various baseline methods are used to evaluate the model’s performance, ranging from random sampling to different SOTA retrieval strategies. The following is a description of the baselines used in our experiments.
@@
[…]
@@
5. DEmO: DEmO is a two‐stage, data‐free framework for configuring an optimal in‐context sequence using influence, a concept that has become increasingly popular in ICD selection. In the first stage, it draws $N$ random permutations $\{\pi_{i}\}_{i=1}^{N}$ of the candidate support set and measures each permutation’s label‐fairness by computing its content‐free entropy:
@@
where $C_{\pi}$ is the prompt constructed by $\pi$ with a content‐free token in place of the query. The top-$K$ permutations with the highest $E(\pi)$ are retained as the candidate set $\Pi$.
@@
Finally, permutation $\pi^{*}=\arg\max_{\pi\in\Pi}I(x_{t};\pi)$ is chosen for the actual prediction. This targeted re‐ranking ensures that each test sample uses the example order most “influential” to its correct classification, without relying on any additional labeled data.
@@ E.4 Results and Analysis
Model Method VQA Captioning Classification Hybrid Fast CLEVR
@@
We can go deep into the per-model results in Table 8. The findings are as follows: (1) TACO exhibits the best performance in all but three tasks across nine datasets and five LVLMs, demonstrating its great efficiency and generalization. Upon examining the outputs, we observe that GPT-4V tends to deviate from the ICD format and produce redundant information more easily than open-source LVLMs, aligning with (Wu et al., 2023a). This results in the quality improvement of the ICL sequence not always translating into stable ICL performance gains for GPT-4V, which may explain why TACO did not achieve the best performance in two of its tasks. (2) For tasks like VizWiz and Hybrid, TACO consistently improves the quality of sequence generation in all LVLMs compared to similarity-based models, demonstrating the importance of increasing task semantics for complex task mappings. We find that the performance gains from TACO are not directly related to the model’s intrinsic ability on these tasks. Unlike simpler tasks like classification, for tasks with complex mappings, task semantics still has a significant impact, even when LVLMs exhibit strong few-shot learning abilities. This shows that models with strong ICL capabilities on certain tasks retain, and even strengthen, their ability to leverage task semantics, underscoring the value of improving ICL sequence quality.
@@
Efficiency analyses. ICL is widely adopted for its efficiency Kang et al. (2025); therefore, efficiency was a primary focus in the design of TACO. TACO exhibits high efficiency during both training and inference. Firstly, TACO is a lightweight language model composed solely of a fusion module and four transformer decoder blocks. With only 140M parameters, its training cost is extremely low compared to LLMs. Moreover, owing to its specialized training objectives, TACO prioritizes data quality over sheer quantity, enabling effective training with a relatively small amount of high-quality data. Simultaneously, our approach to constructing training data is highly efficient. By leveraging LVLM for self-assessment, we significantly reduce the overhead associated with incorporating external metrics. For instance, when using CIDEr to construct training data for the image captioning task, the costs are nearly 9 times higher than those of our current method. To further validate TACO’s training efficiency, we compare its training time with that of 4-layer Lever-LM on the same training sets.
@@ Appendix F Additional Ablation Study
F.1 Input Embeddings
VQAv2 MSCOCO HatefulMemes Hybrid Fast CLEVR
@@ F.1 Input Embeddings
where $f_{I},f_{Q}$ and $f_{R}$ denote the weights computed using a softmax function applied to the linear transformations, ensuring their sum equals 1. Additionally, we apply regularization to the weights: $f_{I}^{2}+f_{Q}^{2}+f_{R}^{2}\leq\theta$ to prevent excessive reliance on specific components.
@@ F.2 Instruction
Table 14: Results of TACO with diverse instruction types. The highest scores are highlighted in bold. $Inst^{\prime}\to Inst$ means using InstInst during the initialization of $TG$.
$^{\prime} InstInst$ VQAv2 VizWiz OK-VQA Hybrid
@@
Table 15: Results of TACO under various $^{\prime}-InstInst$ combinations. $Inst^{\prime}$ represents the style used for initializing TGTG, while InstInst refers to the style actually incorporated into the prompt.
F.3 Generalization Test
@@
For NLP evaluation, we utilize the Rule Learning part of the latest benchmark, ICLEval. ICLEval is designed to assess the ICL abilities of LLMs, focusing on two main sub-abilities: exact copying and rule learning. The Rule Learning part evaluates how well LLMs can derive and apply rules from examples in the context. This includes tasks such as format learning, where models must replicate and adapt formats from given examples, and order and statistics-based rule learning, where the model must discern and implement patterns such as item sequencing or handling duplications. These tasks challenge LLMs to go beyond language fluency, testing their ability to generalize from context in diverse scenarios. Examples of $(Q,R)$ pairs can be found in Table 18. For all tasks, we use exact match scores to evaluate the predictions against the labels.
@@
F.4 Revisiting Task Mapping Framework
@@
Besides evaluating performance, we also employ in-context lens to examine the average evolution of the internal outputs in LVLM by these methods under HM settings. Figure 10 illustrates the evolution of the LVLM’s internal reasoning as it learns from ICL sequences generated by different methods. The results show that sequences produced by TACO enable the model to most effectively infer task mappings and generate accurate responses.